Problem 18-02A (Video Your answer is partially correct. Try again. The comparative statements of Wahlberg Company are presented here. Wahlberg Company ncome Statement For the Years Ended December 31 2020 2019 Net sales Cost of goods solod Gross profit Selling and administrative expenses Income from operations Other expenses and losses $1,816,600 1,010,700 805,900 518,600 287,300 $1,750,100 982,000 768,100 475,600 292,500 Interest expense Income before income taxes Income tax expense Net income 19,000 268,300 81,196 $ 187,104 13,900 278,600 76,900 $ 201,700

anlDerg Compahy Balance Sheets December 31 Assets 2020 2019 Current assets Cash Debt investments (short-term) Accounts receivable Inventory $60,100 68,100 116,200 123,200 367,600 597,700 $965,300 $63,500 49,900 101,100 114,200 328,700 520,800 $849,500 Total current assets Plant assets (net) Total assets Liabilities and Stockholders Equit Current liabilities $145,100 42,900 188,000 200,000 388,000 Accounts payable Income taxes payable $160,300 43,800 204,100 220,000 424,100 Total current liabilities Bonds payable Total liabilities Stockholders equity Common stock ($5 par) Retained earnings 281,300 259,900 541,200 $965,300 303,400 158,100 461,500 $849,500 Total stockholders equity Total liabilities and stockholders equity

All sales were on account. Net cash provided by operating activities for 2020 was $244,000. Capital expenditures were $140,000, and cash dividends were $85,304 Compute the following ratios for 2020. (Round Earnings per share, Current ratio and Asset turnover to 2 decimal places, e.g. 1.65 or 1.65:1, and all other answers to 1 decimal place, eg. 6.8 or 6.8%. Use 365 days for calculation.) (a) Earnings per share 3.33 (b) Return on common stockholders equity c) Return on assets (d) Current ratio (e) Accounts receivable turnover (f) Average collection period (g) Inventory turnover (h) Days in inventory (i) Times interest earned 37.31 % 1.8:1 times times

(i) Asset turnover times (k) Debt to assets ratio (I) Free cash flovw Click if you would like to Show Work for this question: Open Show Work

Help me solve for the red blanks
Answer #1

(a)Earnings per share =Net Income/Weighted average shares outstanding

                shares outstanding         Weights               Weighted average

                60680                                    0.5                          30340

                56260                                    0.5                          28130

                Weighted average shares outstanding   58470

                             =$ 187104/58470=$ 3.2/share

(c)Return on assets= (Net Income*100)/Average assets

   Assets(2019):$849500

Assets (2020):$ 965300

Average assets=($ 849500+$ 965300)/2=$907400

Return on assets = ($ 187104*100)/$ 907400=20.62%

( e)Accounts Receivable Turnover=Net Credit Sales/Average accounts receivable

              =$1816600/($116200+$101100)/2

  Accounts Receivable Turnover =16.7 times

(f)Average Collection Period= 365/Accounts Receivable Turnover

   =365/16.72=21.8 days

(g )Inventory Turnover=times=Cost of goods sold/Average Inventory

                                  =$1010700/($123200+$114200)/2

   Inventory Turnover ratio=$1010700/$118700=8.5 times

(h )Days in Inventory= 365/Inventory Turnover Ratio

                               =365/8.51=42.9 i.e 43 days

(i)Times Interest Earned=Income before interest and taxes/interest expense

                                             =$ 287300/$19000=15.1 times